# Coat stiffening explains the consensus pathway of clathrin-mediated endocytosis

### v1

@@ Abstract
Clathrin-mediated endocytosis is the main pathway used by eukaryotic cells to take up extracellular material, but the dominant physical mechanisms driving this process are still elusive. Recently several high-resolution imaging techniques have been used on different cell lines to measure the geometrical properties of clathrin-coated pits over their whole lifetime. Here we first show that all datasets follow the same consensus pathway, which is well described by the recently introduced cooperative curvature model, which predicts a flat-to-curved transition at finite area, followed by linear growth and subsequent saturation of curvature. We then apply an energetic model for the composite of plasma membrane and clathrin coat to the consensus pathway to show that the dominant mechanism for invagination is coat stiffening, which results from cooperative interactions between the different clathrin molecules and progressively drives the system towards its intrinsic curvature. Our theory predicts that two length scales determine the time course of invagination, namely the patch size at which the flat-to-curved transition occurs and the final pit radius.
@@ I Introduction
The exact time point at which curvature is generated during CME has been debated since many decades [11, 12]. Early work with cells supported the constant area model (CAM), in which clathrin triskelia first grow into large flat patches, which then invaginate. In contrast, reconstitution experiments with lipids and clathrin tend to favor the constant curvature model (CCM), in which curvature is generated right from the start of the clathrin assembly [7]. Today it is accepted that the actual process is a mixture of both scenarios, with both area and curvature growing in time during CME. Recently several high resolution imaging techniques have been used to spatially resolve the time course of CME in mammalian cells [13, 14, 15, 16]. Using electron tomography (ET) [13], high-speed atomic force microscopy (HS-AFM) [17] and super-resolution microscopy (SRM) [16], it has been shown that curvature starts to be generated only once around half of the final clathrin coat has been assembled. What all these techniques have in common is that snapshots are taken of clathrin coats that can be sorted according to the state of invagination, which is used as a surrogate for time. However, until now these different datasets have not been compared with each other. In particular, the HS-AFM data has not yet been evaluated as a function of pseudo-time.
Figure 1: Cooperative curvature model (CoopCM). (A) The clathrin coat is assumed to have the shape of a spherical cap with invagination angle $\theta$, cap radius $R$, area $A$ and edge length $\mathcal{E}$. Curvature $H=1/R$ is therefore the same everywhere, but changes in time, as does the invagination angle $\theta$. (B) Curvature $H$, (C) area $A$ and (D) edge length $\mathcal{E}$ as a function of invagination angle $\theta$ as predicted by the CoopCM, normalized to their maxima (for $H_{0}=\gamma$). The invagination angle $\theta$ acts as surrogate for time.
@@
To provide a simple mathematical description of coat assembly and invagination, we have recently suggested a conceptually transparent and analytically solvable kinetic model, the cooperative curvature model (CoopCM), that gives excellent fits to the growth kinetics observed with SRM [16]. The CoopCM is based on only a few simple assumptions that are motivated by the experimental observations, namely that the growing pit resembles a spherical cap, that area growth occurs mainly at the rim and that curvature is generated everywhere in the lattice, but in a cooperative manner, until a preferred value is reached. Although this phenomenological approach is effective, it does not address the underlying mechanisms. In principle, different physical mechanisms can contribute to drive coat invagination, including generation of spontaneous curvature, coat stiffening and line tension [25, 17].
In this work, we first discuss the CoopCM in more detail and show explicitly why it predicts a flat-to-curved transition at finite coat area. We then combine experimental data from different high-resolution techniques and different mammalian cell lines with the CoopCM to demonstrate that the invagination pathway of clathrin coats follows a universal pathway that we term the consensus pathway. We next combine this universal pathway with an energetic model for the composite of plasma membrane and clathrin coat. We find that CME could not proceed with a constant coat stiffness, as this would involve a prohibitively large energy penalty for the initially flat clathrin coat. Therefore coat stiffening has to occur over time; combined with the spontaneous curvature of the clathrin triskelia, this provides the dominant energetic contribution and drives invagination.
@@ II The cooperative curvature model
Previously, we showed that the kinetics of the invagination of the clathrin coat is well described by the CoopCM [16]. In the following, we briefly summarize the model and the underlying main assumptions. Our main assumption is that the clathrin coat takes the shape of a spherical cap, as observed experimentally [13, 16], which means that it is characterized by only two quantities, namely the invagination angle $\theta$ and the cap radius $R$ (Fig. 1A). The two most important geometrical quantities for kinetic and energetic models are the cap area $A$ and the edge length $\mathcal{E}$, which for the spherical cap are $A=2\pi R^{2}(1-\cos\theta)$ and $\mathcal{E}=2\pi R\sin\theta$, respectively. It is important to note that the formula for the area of the spherical cap includes the flat circular disk in the double limit when $\theta\rightarrow 0$ and at the same time $R\rightarrow\infty$ with the disk radius $R\sin\theta\approx R\theta$ remaining finite.
@@
Instead of giving a fully dynamic description, we describe how the coat curvature $H=1/R$ evolves with $\theta$, which we use as a surrogate of time. The reason for this is that experimentally only snapshots of clathrin coats are taken, which, assuming that coat invagination is an irreversible process, can be ordered according to their opening angle $\theta$. The assumption here is that curvature changes occur over the whole domain of the spherical cap, thus making $H$ the dominant slow dynamic variable, while the coat area is controlled by an independent growth mechanism along the edge of the clathrin coat. Moreover, we assume that as soon as the coat area has reached a certain threshold, coat curvature $H$ builds up with a constant rate $\gamma$, driven by the interactions of the clathrin triskelia, which includes the generation of pentagons.
@@
As the clathrin coat invaginates, the generation of coat curvature eventually has to slow down, because the clathrin coat approaches the final coat curvature $H_{0}=1/R_{0}$, which is set by the curvature of the clathrin triskelia, interactions with other clathrin triskelia and interactions with adaptor and accessory proteins at the membrane [20, 23]. This coat curvature $H_{0}$ is expected to be smaller than the curvature of clathrin cages $H_{\mathrm{c}}=1/R_{\mathrm{c}}$, which is determined solely by the curvature of the clathrin triskelia and the interactions between the other clathrin triskelia [5, 11]. $R_{\mathrm{c}}$ has been measured to be in the range $32.5-50\,\mathrm{nm}$ [7, 26]. For typical pit sizes (typical membrane radius 40 nm [13] plus 15 nm accounting for the thickness of the clathrin coat and its gap to the membrane [27, 7]) the expected value for $R_{0}$ is around $55\,\mathrm{nm}$, which indeed is above the range for radii for the cages.
@@
As the coat reaches the saturation curvature $H_{0}$, the coat curvature will stop increasing and a stable steady state of invagination emerges. Therefore, this curvature acts as a control mechanism for the invagination and effectively sets the pit size. Since the curvature saturation only sets in at the late stage of the invagination process and curvature generation is cooperative, we assume that it is proportional to $H^{2}$ (rather than to $H$, see more discussion below). Together, we now have a simple evolution equation for the coat curvature as a function of the invagination angle
@@
Consequently, the radius of the coat is defined as the inverse of the coat curvature $R(\theta)=1/H(\theta)$, where $R_{0}=1/H_{0}$.
Earlier we have shown that the CoopCM can be formulated also as a fully dynamic description that predicts the characteristic square root dependence of curvature with time that was observed experimentally [16]. Until Section IV, we do not address time explicitly, because we first focus on the geometry of the system, which is sufficient to develop an energetic description. However, we have to clarify at which coat area the invagination process starts. As done earlier for the fully dynamical version of the CoopCM, we now make the assumption that area growth along the edge occurs independently of coat invagination, which includes the possibility that invagination starts at finite area. Therefore, we can solve the area of the spherical cap for coat curvature $H$ and consider the coat curvature as a function of $\theta$ at fixed area $A$. We thus use the area formula to write $H=\sqrt{2\pi(1-\cos\theta)/A}$, which is justified as long as $A>0$. Expanding the coat curvature $H$ for a patch with finite area $A>0$ in $\theta$ around the flat state ($\theta=0$) yields
$H=\sqrt{\frac{\pi}{A}}\left(\theta-\frac{\theta^{3}}{24}\right)+\mathcal{O}(\theta^{5})$ (3)
which is justified as long as $A$ changes on another time scale. To obtain a differential equation for the coat curvature, we assume that the derivate of coat curvature $\mathrm{d}H/\mathrm{d}\theta$ can be written as a polynomial in $H$. Following this idea we expand $\mathrm{d}H/\mathrm{d}\theta$ as a power series up to second order in $H$ to take non-linear effects like cooperativity between clathrin triskelia into account. We get $\mathrm{d}H/{\mathrm{d}\theta}=c_{0}-c_{1}H-c_{2}H^{2}$, where $c_{0},c_{1}$ and $c_{2}$ are expansion coefficients. Now we demand that the ansatz is consistent with Eq. (3). By taking the derivative of Eq. (3), we find that the derivative depends only quadratically on $\theta$. Using Eq. (3) also on the power series we see that $c_{1}=0$, as otherwise a linear term in $\theta$ would appear. Thus internal consistency of the theory implies that the non-linear saturation term of the CoopCM in Eq. (1) follows from the fact that coat invagination starts at a finite $A$, given by Eq. (3). We conclude that the CoopCM in fact predicts a flat-to-curved transition at finite coat area, in agreement with the recent experimental observations [16].
In order to determine the value of the coat area at which invagination starts, we examine the area of the spherical cap using the expression for the coat radius $R(\theta)=1/H(\theta)$ (Eq. (2)) in the limit of a flat disk, i.e., $\theta\rightarrow 0$
@@
As a result, we recover the area of a circular disk. From Eq. (4) we also find a second, geometrical interpretation of $\gamma$ as the inverse of the patch radius $R_{\mathrm{T}}=1/\gamma$ at which the invagination of the flat coat starts. In the limit when the coat approaches maximum invagination, we get
@@
We also note that the CCM is a limiting case of the CoopCM if $\gamma\rightarrow\infty$. The CAM is partially recovered from the CoopCM, when we demand $A(\theta=0)=A(\theta=\pi)$, which is true for $\gamma R_{0}=0.4411$. In this case, the corresponding coat area varies by around $10\%$ between $\theta=0$ and $\theta=\pi$ along the domain. These considerations suggest that $\gamma R_{0}$ can be considered as a measure that indicates how close the CoopCM is to the CCM ($\gamma R_{0}\gg 1$) or to the CAM ($\gamma R_{0}=0.4411$).
Figure 2: Experimental data define a consensus pathway that agrees with the CoopCM. (A-F) Curvature $H$ as a function of invagination angle $\theta$ for different data sets. The rolling mean (red) and the fit of the CoopCM (solid black), the CCM (dashed black) and the CAM (dotted black) are shown. (A) Electron tomography (ET) data [13]. (B) High-speed atomic force microscopy (HS-AFM) data [17]. (C-E) Super-resolution microscopy (SRM) data from different cell lines [16]. (F) Pooled data from all datasets where the ET and SRM data are corrected for curvature. The results of the fits are documented in Tab. 1.
To illustrate the universal pathway of coat invagination as predicted by the CoopCM, in Fig. 1 the curvature (B), area (C) and edge length (D) of the clathrin coat are plotted normalized to their maxima for $H_{0}=\gamma$. Due to the normalization and the fact that we here use $H_{0}=\gamma$, the shape of the curves for curvature, area and edge length are independent of the specific choice of $H_{0}$ and $\gamma$. Apart from the saturation behaviour in curvature, we see that invagination of the clathrin coat starts at a finite initial coat area and therefore also at a finite edge length. The edge length has a maximum in the vicinity of the equator.
@@ III The consensus pathway of CME
We now use the CoopCM, in the form of Eq. (2), to fit experimental data sets of the coat curvature $H$ as a function of the invagination angle $\theta$ that have been recorded using different experimental methods and different cell lines. We use ET data [13], HS-AFM data [17], and SRM data [16]. For the ET data, we converted the measured tip radius of membrane pits as a function of $\theta$ to give the coat curvature $H(\theta)$ assuming the geometry of a spherical cap. Similarly, for the HS-AFM data we converted the measured coat area as a function of the coat radius to give $H(\theta)$. For the SRM data, clathrin-coated pits with negative curvature values and curvature values exceeding a threshold were excluded from the analysis (c3T3: $H>0.014\,\mathrm{nm^{-1}}$, SK-MEL-2: $H>0.016\,\mathrm{nm^{-1}}$, U2OS: $H>0.014\,\mathrm{nm^{-1}}$), similar to what was previously reported (note for U2OS: there was a typo in [16] for the threshold value), because these strongly curved pits lack endocytic marker AP2-GFP and might belong to the Golgi [16]. The HS-AFM data is used as kindly provided by the authors [17]. The different data sets, the rolling means, and the fits according to Eq. (2) are shown in Fig. 2. For the CoopCM, we determine two parameters from the fit, $R_{0}=1/H_{0}$ and $\gamma$, which are both documented in Tab. 1.
@@
Interestingly, the scale of coat curvature shows differences between the different datasets. For SK-MEL-2 cells using ET in Fig. 2A, the curvature values reach around $0.02\,\mathrm{nm^{-1}}$, the largest observed values. For PTK2 cells using HS-AFM in Fig. 2B, we observe intermediate values of coat curvature of around $0.015\,\mathrm{nm^{-1}}$. For c3T3, SK-MEL-2 and U20S cells using SRM in Fig. 2C-E, we find the smallest curvature values of around $0.01\,\mathrm{nm^{-1}}$. Interestingly, the differences between the different cell lines using the same technique in Fig. 2C-E are smaller than the differences between different techniques using the same cell line in Fig. 2A and D. The finding can be corroborated when comparing the different values of the preferred coat radius $R_{0}$ in Tab. 1.
@@
It is natural to assume that this systematic discrepancy is caused by how the clathrin coats are imaged in the different techniques. In the ET data of Avinoam et al. [13], the radii of the membranes invaginated by the clathrin coats are measured. Therefore, we expect ET to give large values of curvature. In the HS-AFM data of Tagiltsev et al. [17], the curvature is measured in unroofed cells directly at the clathrin coat. Therefore, we expect intermediate values of coat curvatures for this techniques. In the SRM data of Mund et al. [16], the clathrin coat is labeled by polyclonal antibodies that bind to both clathrin heavy and light chains of permeabilized cells, with a potential preference to the intracellular side due to a higher density of binding sites. Therefore, we expect small values of coat curvature. These arguments explain why we observe that the values of coat curvature decrease from ET through HS-AFM to SRM (cf. Tab. 1).
@@
Since the kinetics of clathrin coat invagination seem to be very similar for all cell lines and techniques (cf. Fig. 2A-E), we decided to pool the data by combining the values from all measurements. Following our previous reasoning and to exclude any bias in the available data, we corrected the measured data from ET by taking into account the thickness of the clathrin coat and its gap to the membrane. The corrected coat curvature is given by $H=1/(1/H_{\mathrm{pit}}+h_{\mathrm{cc}})$, where $H_{\mathrm{pit}}$ is the measured pit curvature and $h_{\mathrm{cc}}=15\,\mathrm{nm}$ is the estimated thickness correction [7]. Similarly, we corrected the measured data from SRM by taking into account the length of the antibodies. The corrected coat curvature is then given by $H=1/(1/H_{\mathrm{a}}-l_{\mathrm{ca}}$), where $H_{\mathrm{a}}$ is the measured antibody-labeled coat curvature and $l_{\mathrm{ca}}=15\,\mathrm{nm}$ is the estimated length of the clathrin antibody. The curvature-corrected pooled data and the result of the corresponding fit (cf. Tab. 1) then define the consensus invagination pathway for clathrin coated pits (Fig. 2F). We see excellent agreement between experimental data and the CoopCM. We also find that the flat-to-curved transition occurs at around $50\%$ of the final coat area. In the following we analyze the energetics of this consensus pathway.

### v2

# Coat stiffening can explain invagination of clathrin-coated membranes

@@ Abstract
Clathrin-mediated endocytosis is the main pathway used by eukaryotic cells to take up extracellular material, but the dominant physical mechanisms driving this process are still elusive. Recently several high-resolution imaging techniques have been used on different cell lines to measure the geometrical properties of clathrin-coated pits over their whole lifetime. Here we first show that the combination of all datasets with the recently introduced cooperative curvature model defines a consensus pathway, which is characterized by a flat-to-curved transition at finite area, followed by linear growth and subsequent saturation of curvature. We then apply an energetic model for the composite of plasma membrane and clathrin coat to this consensus pathway to show that the dominant mechanism for invagination could be coat stiffening, which might originate from cooperative interactions between the different clathrin molecules and progressively drives the system towards its intrinsic curvature. Our theory predicts that two length scales determine the invagination pathway, namely the patch size at which the flat-to-curved transition occurs and the final pit radius.
@@ I Introduction
The exact time point at which curvature is generated during CME has been debated since many decades [11, 12]. Early work with cells supported the constant area model (CAM), in which clathrin triskelia first grow into large flat patches, which then invaginate. In contrast, reconstitution experiments with lipids and clathrin tend to favor the constant curvature model (CCM), in which curvature is generated right from the start of the clathrin assembly [7]. Today it is accepted that the actual process is a mixture of both scenarios, with both area and curvature growing in time during CME.
Recently different high-resolution imaging techniques have been used to gain insight into the spatiotemporal coordination of CME in mammalian cells [13, 14, 15, 16]. An early study with electron tomography (ET) has demonstrated that the curvature of the clathrin coats is not constant, but increases during invagination [13]. Using high-speed atomic force microscopy (HS-AFM) for imaging and nano-dissection, it has been found that clathrin lattices relax after cutting and thus store elastic energy [17]. With super-resolution microscopy (SRM), it was found that the curvature of the clathrin coat starts to be generated only once around half of the final clathrin coat has been assembled in flat patches and that later it saturates [16]. What all these techniques have in common is that snapshots are taken of clathrin coats that can be sorted according to the state of invagination, where usually the invagination angle is taken as a surrogate for time. However, until now these different datasets have not been compared with each other. In particular, the HS-AFM data has not yet been explicitly evaluated as a function of pseudo-time.
Figure 1: Cooperative curvature model (CoopCM). (A) The clathrin coat is assumed to have the shape of a spherical cap with invagination angle $\theta$, cap radius $R$, area $A$ and edge length $\mathcal{E}$. Curvature $H=1/R$ is therefore the same everywhere, but changes in time, as does the invagination angle $\theta$. (B) Curvature $H$, (C) area $A$ and (D) edge $\mathcal{E}$, as predicted by the CoopCM, plotted against the invagination angle $\theta$ and normalized to their maxima (for $H_{0}=\gamma$). The invagination angle $\theta$ acts as surrogate for time.
@@
To provide a simple mathematical description of coat assembly and invagination, we have recently suggested a conceptually transparent and analytically solvable kinetic model, the cooperative curvature model (CoopCM), that agrees with the growth kinetics observed with SRM [16]. The CoopCM is based on only a few simple assumptions that are motivated by the experimental observations, namely that the growing pit resembles a spherical cap, that area growth occurs mainly at the rim and that curvature is generated everywhere in the lattice, but in a cooperative manner, until a preferred value is reached. Although this phenomenological approach is effective, it does not address the underlying mechanisms. In principle, different physical mechanisms can contribute to drive coat invagination, including generation of spontaneous curvature, coat stiffening and line tension [25, 17].
In this work, we first discuss the CoopCM in more detail and show explicitly why it predicts a flat-to-curved transition at finite coat area. We then combine experimental data from different high-resolution techniques and different mammalian cell lines with the CoopCM to define a consensus pathway of CME. We next combine this consensus pathway with an energetic model for the composite of plasma membrane and clathrin coat. Within the framework of this model, we find that dynamic coat stiffening is a simple way in which invagination can be achieved. Because in the interface model invagination is driven by spontaneous curvature and stiffness penalizes deviations from this spontaneous curvature, high stiffness is required to stabilize the target state, but generates a prohibitively large energy penalty for the initially flat clathrin coat. Therefore, a gradual increase of stiffness could balance these two requirements of the system.
@@ II The cooperative curvature model
Previously, we showed that the kinetics of the invagination of the clathrin coat is described by the phenomenological CoopCM, which is justified as it effectively fits the invagination data of clathrin coats [16]. In the following, we briefly summarize the model and the underlying main assumptions. We also demonstrate that the CoopCM can be derived with both phenomenological and geometrical arguments.
@@
Our first main assumption is that the clathrin coat takes the shape of a spherical cap. This assumption is supported well by electron microscopy [13] as well as by SRM [26, 16] for mammalian cell types (but not for yeast, which in addition has a high turgor pressure). A spherical cap is characterized by only two quantities, namely the invagination angle $\theta$ and the cap radius $R$ (Fig. 1A). The two most important geometrical quantities for kinetic and energetic models are the cap area $A$ and the edge length $\mathcal{E}$, which for the spherical cap read $A=2\pi R^{2}(1-\cos\theta)$ and $\mathcal{E}=2\pi R\sin\theta$, respectively. It is important to note that the formula for the area of the spherical cap includes the flat circular disk in the double limit when $\theta\rightarrow 0$ and at the same time $R\rightarrow\infty$ with the disk radius $R\sin\theta\approx R\theta$ remaining finite.
@@
Instead of giving a fully dynamic description, we describe how the coat curvature $H=1/R$ evolves with $\theta$, which we use as a surrogate of time. The reason for this is that experimentally only snapshots of clathrin coats are taken, which, assuming that coat invagination is an irreversible process, can be ordered according to their opening angle $\theta$. Within the spherical cap assumption, curvature than is the second important variable. Experimentally it has been found that curvature changes the fastest at the beginning of invagination [16]. We therefore assume that curvature $H$ is the dominant fast dynamic variable and that coat area $A$ is controlled by an independent growth mechanism along the edge of the clathrin coat. The initial growth rate is called $\gamma$ and should be determined by the interaction of the clathrin triskelia, which self-assembly into a hexagonal lattice with pentagonal defects.
@@
As the clathrin coat invaginates, the generation of coat curvature has to slow down eventually, because the clathrin coat approaches the final coat curvature $H_{0}=1/R_{0}$, which is set by the curvature of the clathrin triskelia, interactions with other clathrin triskelia and interactions with adaptor and accessory proteins at the membrane [20, 23]. This coat curvature $H_{0}$ is expected to be smaller than the curvature of clathrin cages $H_{\mathrm{c}}=1/R_{\mathrm{c}}$, which is determined solely by the curvature of the clathrin triskelia and the interactions between the other clathrin triskelia [5, 11]. The cage radius $R_{\mathrm{c}}$ has been measured to be in the range $32.5-50\,\mathrm{nm}$ [7, 27]. For typical pit sizes (typical membrane radius 40 nm [13] plus 15 nm accounting for the thickness of the clathrin coat and its gap to the membrane [28, 7]) the expected value for $R_{0}$ is around $55\,\mathrm{nm}$, which indeed is above the range for radii for the cages.
@@
As the coat reaches the saturation curvature $H_{0}$, the coat curvature will stop increasing and a stable steady state of invagination emerges. Therefore, this curvature acts as a control mechanism for the invagination and effectively sets the pit size. Since the curvature saturation only sets in at the late stage of the invagination process, we assume that it is proportional to $H^{2}$. Importantly, the choice of $H^{2}$, rather than of $H$, is also justified phenomenologically, as it gives better fits [16]. Although $H$ is a high-level feature of the system, it is tempting to speculate that the functional form of $H^{2}$ also indicates that curvature saturation is a signature of cooperativity between clathrin triskelia.
Together, we now have a simple evolution equation for the coat curvature as a function of the invagination angle
@@
[…]
Besides this phenomenological derivation, we want to demonstrate that Eq. (1) can be also motivated using geometrical arguments. Earlier we have shown that the CoopCM can be formulated also as a fully dynamic description that predicts the characteristic square root dependence of curvature with time that was observed experimentally [16]. Until Section IV, we do not address time explicitly, because we first focus on the geometry of the system, which is sufficient to develop an energetic description. However, as done earlier for the fully dynamical version of the CoopCM, we now make the assumption that area growth along the edge occurs independently of coat invagination, which includes the possibility that invagination starts at finite area.
Assuming that the clathrin coat takes the shape of a spherical cap, defined by the coat radius $R$ (or its inverse the coat curvature $H$) and opening angle $\theta$, we can solve the area formula for the coat curvature. As a result we obtain $H=\sqrt{2\pi(1-\cos\theta)/A}$, which is justified as long as $A>0$. Expanding the coat curvature $H$ for a patch with finite area ($A>0$) in $\theta$ around the flat state ($\theta=0$) yields
$H=\sqrt{\frac{\pi}{A}}\left(\theta-\frac{\theta^{3}}{24}\right)+\mathcal{O}(\theta^{5})\,.$ (3)
The expansion is justified as long as $A$ changes on another (slower) time scale.
In order to obtain a differential equation for the coat curvature, we assume that the derivative of coat curvature $\mathrm{d}H/\mathrm{d}\theta$ takes the form of a polynomial in $H$. To include non-linear effects like cooperativity between clathrin triskelia we consider the equation to be of second order in $H$, $\mathrm{d}H/{\mathrm{d}\theta}=c_{0}-c_{1}H-c_{2}H^{2}$, where $c_{0},c_{1}$ and $c_{2}$ are expansion coefficients. We next use the expression for $H$ of Eq. (3) on the polynomial and compare the resulting equation with the derivative of Eq. (3) with respect to $\theta$. From the comparison of the two equations, we find that $c_{1}=0$ since the derivative of Eq. (3) does not include a linear term in $\theta$. Thus internal consistency of the theory implies that the non-linear saturation term of the CoopCM in Eq. (1) follows from the fact that coat invagination starts at a finite $A$, given by Eq. (3). We conclude that the CoopCM in fact predicts a flat-to-curved transition at finite coat area, in agreement with recent experimental observations [16]. Although the geometric motivation is consistent, we note that different microscopic models could lead to the same macroscopic description.
We still have to clarify at which coat area the invagination process starts. We therefore examine the area of the spherical cap using the expression for the coat radius $R(\theta)=1/H(\theta)$ (Eq. (2)) in the limit of a flat disk, i.e., $\theta\rightarrow 0$
@@
We see that the CoopCM indeed automatically takes care of the double limit required to get a flat disc as initial condition. From Eq. (4) we also find a second, geometrical interpretation of $\gamma$ as the inverse of the patch radius $R_{\mathrm{T}}=1/\gamma$ at which the invagination of the flat coat starts. In the limit when the coat approaches maximum invagination, we get
@@
We also note that the CCM is a limiting case of the CoopCM if $\gamma\rightarrow\infty$. The CAM is partially recovered from the CoopCM, when we demand $A(\theta=0)=A(\theta=\pi)$, which is true for $\gamma R_{0}=0.4411$. In this case, the corresponding coat area varies by around $10\%$ between $\theta=0$ and $\theta=\pi$ along the domain. These considerations suggest that the product $\gamma R_{0}$ can be taken as a measure which indicates how similar the CoopCM is with respect to the CCM ($\gamma R_{0}\gg 1$) and CAM ($\gamma R_{0}=0.4411$).
Figure 2: Experimental data define a consensus pathway that agrees with the CoopCM. (A-H) Curvature $H$ as a function of invagination angle $\theta$ for different data sets. The rolling mean (red) and the fit of the CoopCM (solid black), the CCM (dashed black) and the CAM (dotted black) are shown. (A) Electron tomography (ET) data [13]. (B) High-speed atomic force microscopy (HS-AFM) data [17]. (C-E) Super-resolution microscopy (SRM) data from different cell lines [16]. (F) Datasets from A-E and the corresponding individual fits of the CoopCM. (G) Datasets from A-E that are corrected for curvature together with the individual fits of the CoopCM. (H) Pooled data from the datasets A-E, with the ET and SRM data corrected for curvature. The results of the fits are documented in Tab. 1.
To illustrate the universal pathway of coat invagination as predicted by the CoopCM, in Fig. 1 the curvature (B), area (C) and edge length (D) of the clathrin coat are plotted, normalized to their maxima for $H_{0}=\gamma$. Due to the normalization and the fact that we here use $H_{0}=\gamma$, the shape of the curves for curvature, area and edge length are independent of the specific choice of $H_{0}$ and $\gamma$. Apart from the saturation behaviour in curvature, we see that invagination of the clathrin coat starts at a finite initial coat area and therefore also at a finite edge length. The edge length has a maximum in the vicinity of the equator.
@@ III The consensus pathway of CME
We now use the CoopCM, in the form of Eq. (2), to fit experimental data sets of the coat curvature $H$ as a function of the invagination angle $\theta$ that have been recorded using different experimental methods and different cell lines. We use ET data [13], HS-AFM data [17], and SRM data [16]. The HS-AFM data is used as kindly provided by the authors [17].
For the ET data, we converted the measured tip radius of membrane pits as a function of $\theta$ to get the coat curvature $H(\theta)$ assuming the geometry of a spherical cap. Similarly, for the HS-AFM data we converted the measured coat area as a function of the coat radius to obtain $H(\theta)$. For the SRM data, clathrin-coated pits with negative curvature values and curvature values exceeding a threshold were excluded from the analysis (c3T3: $H>0.014\,\mathrm{nm^{-1}}$, SK-MEL-2: $H>0.016\,\mathrm{nm^{-1}}$, U2OS: $H>0.014\,\mathrm{nm^{-1}}$), similar to what was previously reported (note for U2OS: there was a typo in [16] for the threshold value), because these strongly curved pits lack endocytic marker AP2-GFP and might belong to the Golgi [16].
@@
When we plot the datasets within a single panel, we find that all of them follow similar trends (Fig. 2F). But we also find characteristic differences. Most importantly, the scale of coat curvature is different for the various datasets. For SK-MEL-2 cells using ET in Fig. 2A, the curvature values reach around $0.02\,\mathrm{nm^{-1}}$, the largest observed values. For PTK2 cells using HS-AFM in Fig. 2B, we observe intermediate values of coat curvature of around $0.015\,\mathrm{nm^{-1}}$. For c3T3, SK-MEL-2 and U20S cells using SRM in Fig. 2C-E, we find the smallest curvature values of around $0.01\,\mathrm{nm^{-1}}$. Interestingly, the differences between the different cell lines using the same technique in Figs. 2C-E are smaller than the differences between different techniques using the same cell line in Figs. 2A and D. The finding can be corroborated when comparing the different values of the preferred coat radius $R_{0}$ in Tab. 1.
@@
It is natural to assume that this systematic discrepancy is caused by how the clathrin coats are imaged in the different techniques. In the ET data of Avinoam et al. [13], the radii of the membranes invaginated by the clathrin coats are measured. We therefore expect ET to give large values of curvature. In the HS-AFM data of Tagiltsev et al. [17], the curvature is measured in unroofed cells directly at the clathrin coat. We thus expect intermediate values of coat curvatures for this techniques. In the SRM data of Mund et al. [16], the clathrin coat is labeled by polyclonal antibodies that bind to both clathrin heavy and light chains of permeabilized cells, with a potential preference to the intracellular side due to a higher density of binding sites. Therefore, we expect small values of coat curvature. These arguments explain why we observe that the values of coat curvature decrease from ET through HS-AFM to SRM (cf. Tab. 1).
@@
Since the kinetics of clathrin coat invagination seem to be very similar for all cell lines and techniques (cf. Figs. 2A-F), we decided to pool the data by combining the values from all measurements. Following our previous reasoning and to exclude any bias in the available data, we corrected the measured data from ET by taking into account the thickness of the clathrin coat and its gap to the membrane. The corrected coat curvature is given by $H=1/(1/H_{\mathrm{pit}}+h_{\mathrm{cc}})$, where $H_{\mathrm{pit}}$ is the measured pit curvature and $h_{\mathrm{cc}}=15\,\mathrm{nm}$ is the estimated thickness correction [7]. Similarly, we corrected the measured data from SRM by taking into account the length of the antibodies. The corrected coat curvature is then given by $H=1/(1/H_{\mathrm{a}}-l_{\mathrm{ca}}$), where $H_{\mathrm{a}}$ is the measured antibody-labeled coat curvature and $l_{\mathrm{ca}}=15\,\mathrm{nm}$ is the estimated length of the clathrin antibody.
For the curvature-corrected data, the individual data sets are fitted in Fig. 2G. The scattering in the experimental data is considerable, which might be partially due to resolution limits. Moreover, since CME in cells involves many different types of proteins that are all underlying stochastic fluctuations, CME itself is highly stochastic. In addition, there are abortive events and there might be subpopulations, both of which are not taken into account here. However, we also note that the variability within single data sets is larger than the variability between the fits of the individual data sets, which seems to indicate that CME is rather similar across imaging methods and cell types. Thus, we decided to fit and evaluate the pooled data in order to find the mean trajectory for CME (cf. Fig. 2H and Tab. 1). We see good agreement between experimental data and the CoopCM and conclude that the resulting consensus pathway can be used as a reasonable basis for an energetic analysis. We also find that the flat-to-curved transition occurs at around $50\%$ of the final coat area.